OpenStudy (anonymous):

HELPHELPHELP!!!!!!!!!!1

OpenStudy (anonymous):

Identify the vertex for the graph of y = −3x2 + 6x − 4.

OpenStudy (compassionate):

Use the vertex formula. \[-\frac{ b }{ 2a }\]

OpenStudy (anonymous):

i don't get how to use it?

OpenStudy (compassionate):

a -3 b = 6 Now plug those numbers in and simplify.

OpenStudy (anonymous):

-1?

OpenStudy (compassionate):

\[-\frac{ 6 }{ 2(-3) } \rightarrow \frac{ -6 }{ -6} \rightarrow ?\]

OpenStudy (anonymous):

so its 1?

OpenStudy (compassionate):

It's POSITIVE 1. Yes.

OpenStudy (anonymous):

what abput the second number?

OpenStudy (compassionate):

Plug your 1 into the equation

OpenStudy (anonymous):

5? @compassionate

OpenStudy (compassionate):

Yes